# After Retrieval, Before Generation: Enhancing the␣Trustworthiness of Large Language Models in RAG

### v1

@@ B.4 All System Prompts in BRIDGE
Our framework employs distinct response strategies based on the predicted decision: (1) For FA, the LLM generates responses using all available knowledge sources ($K_{int}$, $K_{gen}$, $K_{ext}$, $K_{ret}$) to provide comprehensive answers. (2) For FI, the LLM relies solely on internal knowledge ($K_{int}$) and generated knowledge ($K_{gen}$), ensuring factual consistency without external noise. (3) For FE, the response is derived from external knowledge ($K_{ext}$) and retrieved temporal knowledge ($K_{ret}$), prioritizing up-to-date information. (4) For RA, the LLM responds with "I don’t know" to avoid propagating potentially toxic or misleading content from adversarial external knowledge.
@@
Reflection Module ==Input data== Original question: {question} Knowledge document: Internal knowledge: {internal_knowledge} External knowledge: {external_knowledge} Generated knowledge: {generated_knowledge} Retrieved knowledge: {retrieved_knowledge} Phase 1: Knowledge contradiction analysis Please perform the following analysis steps: Consistency verification: 1. Confirm the consistency performance of internal knowledge and generated knowledge. 2. Confirm the consistency performance of external knowledge and retrieved knowledge. 3. Mark the specific contradictions between internal and external knowledge. Contradiction classification (analyzed from the following dimensions): Factual contradiction (objective fact difference) Timeliness contradiction (new and old information difference) Perspective contradiction (position/viewpoint difference) Integrity contradiction (information coverage difference) Root cause analysis: Model knowledge limitation (training data/time cutoff) External knowledge bias (source reliability/update frequency) Retrieval matching error (query-document relevance) Generate hallucination problem Phase 2: Problem reconstruction requirements Based on the above analysis, please design number new wildly diverse questions with different words that have the same answer as original question. Requirements: 1. Use different sentence structures. 2. Each question must employ a unique interrogative word (how/which/why, etc.). 3. Cover multiple dimensions of problem-solving. 4. Finally, rank the questions in descending order of importance for each dimension. 5. Pay attention to checking knowledge contradictions in the questions. Output format [Knowledge contradiction analysis] [Main contradiction] [Contradiction type] [Possible cause] [Reconstruct question list] (in descending order of importance) New questions: 1. [New Question 1] 2. [New Question 2] {number}. [New Question {number}]
B.5 Case Study
As shown in Figure 8, in the RA scenario, BRIDGE encounters a situation where the external knowledge is incorrect while the model’s internal knowledge remains empty. After retrieval and generation, the model fails to acquire valid knowledge, resulting in consistently low scores that lead to classification as a rejection scenario. In contrast, vanilla RAG methods are susceptible to interference from erroneous external knowledge, causing them to incorrectly select options like "B.Dream Works Animation" This demonstrates BRIDGE’s capability in handling knowledge-deficient scenarios through its rejection strategy.
@@ B.5 Case Study
Appendix C Baseline Settings
We segment the Wikipedia dump into passages of 256 tokens using sentence boundaries as delimiters, with no overlapping segments, and encode them using BGE-m3. This dump is contaminated with noisy data from the TRD dataset to simulate real-world scenarios involving unreliable external knowledge. For BRIDGE, the retriever defaults to fetching the single most relevant passage per sub-query. For other baselines, we allow retrieving up to 10 passages per question. This typically provides baselines with more retrieved knowledge than BRIDGE, representing a potential advantage. To ensure fair comparison, all models are given access to the same complete set of knowledge sources. Under ideal conditions, this setup ensures that their performance is not attributable to knowledge source availability constraints. We selected state-of-the-art representative baselines for evaluating trustworthiness in the RAG system. For the USC method, we permit 5 API calls to generate 5 responses while allowing integration of both internal and external knowledge sources. RobustRAG is configured to invoke an API for summarizing each retrieved passage, followed by keyword aggregation, resulting in over 10 API calls per question. For AstuteRAG and TrustRAG, whose performance typically correlates with knowledge integration iterations, we adopt the maximum values reported in their papers, ultimately requiring 4 API calls.
As shown in Table 10, we compare the API cost and efficiency between our method and existing approaches. To eliminate internal knowledge gaps across different LLM families, we directly generate the internal knowledge. Consequently, these generations are not counted toward the total API calls. As a result, BRIDGE requires an optimal of 4 API calls in the ideal case and a maximum of 7 in the worst case. While our method incurs slightly more API calls than some knowledge integration baselines, it achieves significantly better overall performance. We believe this represents a favorable trade-off, as the marginal increase in computational cost yields substantial gains in capability.
Table 10: API efficiency.
TRD ConflictBank
Methods Avg API Call Acc Efficiency Avg API Call Acc Efficiency
USC 6 65.53 10.87 6 73.46 12.24
RobustRAG 11 62.50 5.68 11 58.40 5.31
AstuteRAG 4 64.26 16.07 4 47.45 11.86
TrustRAG 4 66.39 16.60 4 30.00 7.50
BRIDGEICL 4.34 72.32 16.66 4.21 61.86 14.69
BRIDGEGRPO 4.22 77.90 18.46 4.14 78.12 18.87
Regarding fine-tuned approaches (SelfRAG, CAD, InstructRAGFT, and Trust-Align), we reproduce their training procedures by implementing the reported hyper-parameters and source code to train all models on Llama3-8B-Instruct. Notably, we adapt SelfRAG (originally implemented on Llama2-7B) following InstructRAGFT’s reproduction method, employing a learning rate of 1e-5 for 2 epochs. All models receive identical computational resources during training and inference phases.

### v2

@@ B.3 All System Prompts in BRIDGE
Appendix C Experiment Settings
C.1 RAG Settings
We segment the Wikipedia dump into non-overlapping passages of 256 tokens using sentence boundaries as delimiters and then encode them with BGE-m3. This dump is contaminated with noisy data from the TRD dataset to simulate real-world scenarios involving unreliable external knowledge. For BRIDGE, the retriever defaults to fetching the single most relevant passage per sub-query. For other baselines, we allow retrieving up to 10 passages per question. This typically provides baselines with more retrieved knowledge than BRIDGE, representing a potential advantage. To ensure a fair comparison, all models are given access to the same complete set of knowledge sources. Under ideal conditions, this setup ensures that their performance is not attributable to knowledge source availability constraints.
C.2 Baseline Settings
We select representative state-of-the-art baselines for evaluating trustworthiness in RAG systems. For the USC method, we permit 5 API calls to generate 5 responses while allowing integration of both internal and external knowledge sources. RobustRAG is configured to invoke an API for summarizing each retrieved passage, followed by keyword aggregation, resulting in over 10 API calls per question. For AstuteRAG and TrustRAG, whose performance typically improves with more knowledge integration iterations, we adopt the maximum iteration counts reported in their respective papers, resulting in 5 API calls per question.
For the fine-tuned approaches (SelfRAG, CAD, InstructRAGFT, and Trust-Align), we reproduce their training procedures using the reported hyper-parameters and source code, training all models on the Llama3-8B-Instruct base. Notably, we adapt SelfRAG (originally implemented on Llama2-7B) following InstructRAGFT’s reproduction method, employing a learning rate of 1e-5 for 2 epochs. All models receive identical computational resources during training and inference phases.
For a fair comparison, all methods employ identical retriever parameters and access the same Wikipedia dumps (cutoff date: February 1, 2025). Furthermore, the hyperparameters for the BRIDGE framework, specifically its weights and thresholds, are determined via grid search on the TRD dataset and then directly deployed on all other datasets without further tuning.
Appendix D Reflection Module
Figure 7: Acc under different reflection iterations.
In our experiments, the reflection mechanism is activated in approximately 10% of all cases. As shown in Figure 7, within these activated instances, the model’s problem-solving accuracy positively correlates with the number of reflection iterations. However, as a plug-and-play component, we carefully considered the efficiency trade-offs. Under ideal conditions, BRIDGE can resolve questions with 5 LLM API calls. When accounting for reflection operations, the average number of API calls per query reaches 5.34 on the TRD dataset - a number comparable to other prompt-based methods. Importantly, this modest increase in computational overhead enables BRIDGE to maintain balanced performance across diverse scenarios. The system prompt is as follows:
@@ Appendix D Reflection Module
Reflection Module Prompt ==Input data== Original question: {question} Knowledge document: Internal knowledge: {internal_knowledge} External knowledge: {external_knowledge} Generated knowledge: {generated_knowledge} Retrieved knowledge: {retrieved_knowledge} Phase 1: Knowledge contradiction analysis Please perform the following analysis steps: Consistency verification: 1. Confirm the consistency performance of internal knowledge and generated knowledge. 2. Confirm the consistency performance of external knowledge and retrieved knowledge. 3. Mark the specific contradictions between internal and external knowledge. Contradiction classification (analyzed from the following dimensions): Factual contradiction (objective fact difference) Timeliness contradiction (new and old information difference) Perspective contradiction (position/viewpoint difference) Integrity contradiction (information coverage difference) Root cause analysis: Model knowledge limitation (training data/time cutoff) External knowledge bias (source reliability/update frequency) Retrieval matching error (query-document relevance) Generate hallucination problem Phase 2: Problem reconstruction requirements Based on the above analysis, please design number new wildly diverse questions with different words that have the same answer as original question. Requirements: 1. Use different sentence structures. 2. Each question must employ a unique interrogative word (how/which/why, etc.). 3. Cover multiple dimensions of problem-solving. 4. Finally, rank the questions in descending order of importance for each dimension. 5. Pay attention to checking knowledge contradictions in the questions. Output format [Knowledge contradiction analysis] [Main contradiction] [Contradiction type] [Possible cause] [Reconstruct question list] (in descending order of importance) New questions: 1. [New Question 1] 2. [New Question 2] {number}. [New Question {number}]
@@
Appendix E Case Study
As shown in Figure 8, in the RA scenario, BRIDGE encounters a situation where the external knowledge is incorrect while the model’s internal knowledge remains empty. After retrieval and generation, the model fails to acquire valid knowledge, resulting in consistently low scores that lead to classification as a rejection scenario. In contrast, vanilla RAG methods are susceptible to interference from erroneous external knowledge, causing them to incorrectly select options like "B.Dream Works Animation". This demonstrates BRIDGE’s capability in handling knowledge-deficient scenarios through its rejection strategy.
Appendix F Related Works
Retrieval-augmented Generation. The paradigm of RAG has evolved significantly since its inception, driven by the need to build LLMs on external knowledge while mitigating hallucinations. Early foundational work established the core framework of RAG Guu et al. (2020); Lewis et al. (2020); Borgeaud et al. (2022). Subsequently, its retrieval module has been optimized through techniques such as query rewriting Zheng et al. (2023); Dai et al. (2023); Chan et al. (2024), passage reranking Glass et al. (2022); Yu et al. (2024), and adaptive efficient retrieval strategies Asai et al. (2023); Jiang et al. (2023); Yao et al. (2023). In the generation module, recent efforts have focused on improving generation efficiency via prompt compression Jiang et al. (2024); Xu et al. (2024); Cheng et al. (2024) and decoding optimization Merth et al. (2024); Jin et al. (2024a); Wang et al. (2025). These approaches primarily enhance individual RAG components in non-adversarial settings, whereas our research focuses on adversarial conditions (internal and external knowledge containing noise).
Trustworthiness of LLMs in RAG. Ensuring the trustworthiness of LLMs in RAG systems has become a critical research area, particularly in adversarial environments. In scenarios where internal and external knowledge sources conflict, some studies align model responses with internal knowledge Pan et al. (2023); Xu et al. (2023); Chen et al. (2024c); Hong et al. (2024); Weller et al. (2024), while others prioritize external knowledge Li et al. (2023); Gekhman et al. (2023); Zhou et al. (2023). Beyond these binary choices, recent efforts to improve the trustworthiness of LLM responses attempt to merge conflicting sources by using contrastive decoding Jin et al. (2024b); Shi et al. (2024) or by integrating consistent portions of internal and external knowledge Xiang et al. (2024); Wang et al. (2024); Wei et al. (2025); Zhou et al. (2025). However, these methods rely on carefully designed prompt templates to assess knowledge consistency. This matching strategy is sensitive to textual variations and lacks mechanisms for handling unanswerable questions. Previous work on enhancing LLM refusal capabilities in RAG systems relies on building expert knowledge bases Cao (2024) or fine-tuning generative models Song et al. (2025). However, these methods often fail to integrate multiple knowledge sources and do not provide plug-and-play compatibility with black-box, model-based RAG systems. Collectively, these works are orthogonal, typically targeting only one or two scenarios.